OpenStudy (zepp):

Let \(k=\sqrt{n+\sqrt{n+\sqrt{n}+\sqrt{n}+\sqrt{n}+ ~...}}\), \(n \in \mathbb{N}\) Show that \(k\) only tends to a whole number only if \(n\) is the product of two consecutif whole number. This square root thing is troubling me, it goes to infinity, how can I express it under a kind of series, do I have to use the telescopic method?

OpenStudy (zepp):

@TuringTest

OpenStudy (anonymous):

just notice \[k^2=n+\sqrt{n+\sqrt{n+\sqrt{n+\sqrt{n}+...}}}\]

OpenStudy (zepp):

So basically I send all n and square root to k's side, expand and simplify?

OpenStudy (anonymous):

sorry my english is not well

OpenStudy (anonymous):

what is the equivalent of second term in the right hand of equation i posted for u?

OpenStudy (zepp):

:| No idea D:

OpenStudy (anonymous):

\[k^2=n+\sqrt{n+\sqrt{n+\sqrt{n+\sqrt{n}+...}}}=n+k\]

OpenStudy (zepp):

\(n+k=k^2\)? o.o that doesn't make sense D:

OpenStudy (anonymous):

then n=k(k-1) i think that is what u want

OpenStudy (zepp):

Where does n=k(k-1) come from?

OpenStudy (zepp):

Ohh, ok ok

OpenStudy (zepp):

http://puu.sh/Ejof Like this?

OpenStudy (anonymous):

thats right

OpenStudy (zepp):

So k and k-1 are two consecutive whole numbers, n is the product of two consecutive whole numbers and a natural number, therefore k is a whole number?

OpenStudy (anonymous):

yes

OpenStudy (zepp):

Great! Thank you so much! :D

OpenStudy (anonymous):

your welcome